A trolley with mass m1 of 3.3 kg is connected to a light string with a weight of mass m2 of 2.2 kg.
The cord runs through a pulley (a small wheel where the cord runs freely)
a) Calculate the vehicle acceleration
b) Calculate the clamping force in the card
1
Expert's answer
2021-07-16T10:21:10-0400

Mass of trolley "(m_1)=3.3 kg"

Mass of connected weight "(m_2)=2.2kg"

Now, Let the common acceleration be a and the clamping force be T (tension in the string).

Now,

"T=m_1a....(i)"

"m_2g-T=m_2a...(ii)"

From equation (i) and (ii)

"a=\frac{m_2g}{m_1+m_2}"

Now, substituting the values,


"\Rightarrow a =\frac{2.2\times 10}{3.3+2.2}m/s^2"


"\Rightarrow a = \frac{22}{5.5}m/s^2"


"\Rightarrow a = 4m/s^2"

hence, the required tension in the string or the clamping force

"T=m_1a = 3.3\times 4 N"

"\Rightarrow T= 13.2N"